24. Which structure for dinitrogen sulfide (N-N S or NSN) is the more stable? a. N N S, because the formal charges can all be zero. b. N-NS, because the formal charges can be reduced to the lowest possible values I charges can be reduced to the lowest possible values d. N-S-N, because the formal charges can all be zero. s electronegative than nitrogen.
Answer #1

24)

Formal charge = ( Number of valence electrons on atom) -(non-bonding valence electrons+ bonding electrons/2)

For N-N-S

Formal charge: 5-(2+6/2) = (0) , 5-(0+8/2) = (+1) , 6-(6+2/2)= (-1) for N, N, S respectively

We have to chose the structure whose formal charges closest to zero i.e as low as possible

option B is the right answer.